OpenStudy (lonnie2018):

put y + 4 = – 1/3 (x – 6) in slope-intercept form?

OpenStudy (shadowlegendx):

For this, all you have to do is use Algebra 1 skills to solve. Remember that y intercept form is y = mx + b where y is any given y point m is the slope x is any given x point and b is the y intercept y + 4 = – 1/3 (x – 6) First, distribute -1/3 to x and -6. y + 4 = -1/3x + 2 Now subtract 4 from both sides to isolate y and you should get y = -1/3x -2